ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

14. Mathematical Aspects of Computer Science

1991 MS Classification: 49-02 Das, Laxminarayan, Deptt. of Comp. Sc., Texas A & M University, USA Convexity to Invexity Now, Convex sets and functions are not confined to the study of abstract mathematical problems, rather those are extensively used in solving some real world problems. However, many real world problems don’t have a convexity affinity. To avoid those difficulties, the researchers some times modify these concepts. As a result these concepts have grown to a large family obtaining a vast literature. It is essential to keep a track of the subsequent addition. In this poster, we like to present various convex sets, abstract convex sets, convex functions, generalized convex functions, their subsequent developments, application to problem solving and relationship with some nonconvex functions.